OpenStudy (anonymous):

If (c/x) + 1 = b & b doesn't equal 1 then what is x?

OpenStudy (anonymous):

\[\frac{c}{x}+1=b\] \[\frac{c}{x}=b-1\] \[\frac{x}{c}=\frac{1}{b-1}\] \[x=\frac{c}{b-1}\]

OpenStudy (anonymous):

thank you.<3 That's what I got, just making sure it was right, thank youu!

OpenStudy (anonymous):

yw